OpenStudy (anonymous):

This is involving the equation of a line through a given point, parallel to a second line. Assume that the line passes through the point (-3,-2) and is parallel to the line through the points (5,-1) and (-2,-5). Find the equation of this line in slope intercept form. y=

OpenStudy (anonymous):

Slope of line through (5,-1) and (-2,-5) = (-5+1)/(-2-5) = -4/-7 = 4/7 So, parallel line through -3,-2 will have same slope. y = 4x/7 + b -2 = 4(-3)/7 + b b = -2 + 12/7 b = (-14+12)/7 b = -2/7 y = 4x/7 - 2/7 => 7y = 4x - 2

Directrix (directrix):

In slope-intercept form, y = (4/7) x - (2/7)

OpenStudy (anonymous):

I have a word problem you might try